If you want to discharge a 8.00 μF capacitor to 55.0% of its initial charge in 9.22 ms, what value resistor will you need?

Question 2 If you want to discharge a 8.00 μ F capacitor to 55.0% of its initial charge in 9.22 ms, what value resistor will you need? Answer:
Answer #1

HIRO t- RC F- c In(o,55 3 WCo 尺 1927-8レ
Answer #2

SOLUTION :


We use the following formula :


Q = Q0 e^(- t / R C) 

=> Q / Q0 = e^ ( - t / RC)


where, 


Q / Q0 = 0.55 (capacitor is discharged to 55% of initial value)

t = time taken for discharge to 0.20 Q0  is 9.22 ms = 9.22^10^(- 3) sec.

R = Resistance  in Ωs  to be determined .

C = Capacitor capacitance = 8.0 µF= 3.0 *10^(-6) F .


Hence,


0.55 =  e^(- 9.22*10^(-3) / (R * 8.0 *10^(-6)) 

=> 0.55 = e^(- 1152.5 / R)

Taking natural logarithm :

=> ln(0.55) = - 1152.5 / R

=> R = - 1152.5 / ln(0.55) 

=> R = 1927.78   approx. 


=> Resistor of 1927.78  Ω approx. (ANSWER).

answered by: Tulsiram Garg